# Evolver: Chain-of-Evolution Prompting to Boost Large Multimodal Models for Hateful Meme Detection

### v1

@@ 4.1. More Experiment Details
Table 4. Summary of hateful meme detection datasets.
4.2. Datasets, Metrics, and Settings
Datasets    We evaluate Evolver on three widely-used public datasets, namely, Facebook Hateful Meme dataset (FHM) (Kiela et al., 2020), Harmful Meme dataset (HarM) (Pramanick et al., 2021), and Multimedia Automatic Misogyny Identification (MAMI) (fersini-etal-2022-semeval).
Settings    We implement our Evolver based on MMICL (Zhao et al., 2023) and LLaVA-1.5 (Liu et al., 2023a). More details refer to the Appendix 4.1.
4.3. Quantitative Analysis
Table 2 shows the results of our method on three hateful meme detection datasets under the zero-shot setting. As we can see, our method consistently outperforms open-source LMM-based models across all datasets and metrics, which are closely aligned with API-based methods. Compared with the vanilla MMICL, our Evolver with the same architecture achieves higher accuracy (63.6% v.s. 60.4%) on the FHM dataset.
4.4. Ablation Study
4.5. Qualitative Analysis
We visualize the paired evolution memes retrieved by the evolutionary pair mining modules in Figure 2, showing how the evolution information influences the final prediction. The origin model (MMICL) without considering the meme’s evolution, predicts non-hateful because it is not easy to detect the boy’s disability or genetic conditions. After obtaining the evolution information like “discriminates against individuals with disabilities or genetic conditions.”, our Evolver rectify the prediction, which supports the rationale of our method.
4.6. Computing Resource Requirements
@@
4.7. Impact of the Number of the Evolutionary Memes
Figure 3. Effect of the number of evolutionary memes.
5. Additional Case Analysis
To show the effectiveness of our method and align a better understanding of the hatefulness in the dataset, we show more examples generated by Evolver and the baseline model in Fig. 4.
Figure 4. Example results of the Evolver (Ours) and the baseline model (MMICL).
6. Conclusion
We present Evolver to seamlessly boost LMMs for hateful meme detection via Chain-of-Evolution prompting. By integrating the evolution of memes, our method can adapt to unseen memes. Experimental results demonstrate the effectiveness of Evolver.
Impact Statements
Ethics Statement.    Our Evolver aims to detect hateful memes by leveraging large multimodal models (LMMs) through a chain-of-evolution promotion strategy. While our objective is to mitigate the spread and impact of hate speech online, we recognize the potential for misuse or unintended consequences of this technology. Specifically, there are concerns about privacy, censorship, and the potential for biases within the AI model that could result in unfair or inaccurate content moderation. We are committed to ensuring that our technology is used ethically and responsibly.

### v2

@@ Appendix B Data Details
Table 4: Summary of hateful meme detection datasets.
FHM Kiela et al. (2020) curate a dataset of various hateful memes collected on the web to help build models to detect hateful memes. As shown in Table 4, it has 8,500 training examples and 1,000 test examples. The definition of a hateful meme is:
"A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. We define attack as violent or dehumanizing (comparing people to non-human things, e.g. animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hate speech."
Table 5: The definition of hatefulness in the FHM.
MAMI Fersini et al. (2022) propose a dataset of Misogyny memes, which contains 9,948 training examples as well as 1,000 test examples. The definition of Misogyny given by MAMI is:
"meme is misogynous if it conceptually describes an offensive, sexist, or hateful scene (weak or strong, implicitly or explicitly) having as target a woman or a group of women. Misogyny can be expressed in the form of shaming, stereotype, objectification, and/or violence."
Table 6: The definition of hatefulness in the MAMI.
HarM Pramanick et al. (2021) builds a dataset of harmful meme related to Covid. As shown in Table 4, this dataset has 3,013 training samples and 354 test examples. Additionally, HarM gives the definition of a harmful meme as follows:
"Multi-modal unit consisting of an image and an embedded text that has the potential to cause harm to an individual, an organization, a community, or society"
Table 7: The definition of hatefulness in the HarM.
Appendix C Computing Resource Requirements
@@
Appendix D Prompt Engineering
In this section, we show the detailed prompt of Evolution Information Extractor (EIE) for MAMI and HarM datasets in Table 8 and Table 9.
Extract the common misogynous feature of these image caption pairs based on the following misogynous rules: meme is misogynous if it conceptually describes an offensive, sexist, or hateful scene (weak or strong, implicitly or explicitly) having as target a woman or a group of women. Misogyny can be expressed in the form of shaming, stereotype, objectification, and/or violence.
Table 8: The prompt design for EIE in the MAMI.
Extract the common harmful feature of these image caption pairs based on the following misogynous rules: meme is harmful if Multi-modal unit consisting of an image and an embedded text that has the potential to cause harm to an individual, an organization, a community, or society.
Table 9: The prompt design for EIE in the HarM.
Furthermore, we show the detailed prompt of our Evolver in the final prediction for MAMI and HarM datasets in Table 10 and Table 11.
Determine if an image <image0> with its caption: $ocr_{text}$ is misogynous or not misogynous. Below is the misogynous analysis of 5 evolutional image text pairs, where our target image text pair is derived from those evolution image text pairs. Use the misogynous of evolution to help you make decision. Evolution: <Info> Requirement:
1. The output should be a word of either ’misogynous’ or ’not misogynous’. 2. meme is misogynous if it conceptually describes an offensive, sexist, or hateful scene (weak or strong, implicitly or explicitly) having as target a woman or a group of women. Misogyny can be expressed in the form of shaming, stereotype, objectification, and/or violence. [Here is your expert response]
Table 10: Prompt design for final prediction of MAMI.
Determine if an image <image0> with its caption: $ocr_{text}$ is harmful or not harmful. Below is the harmfulness analysis of 5 evolutional image text pairs, where our target image text pair is derived from those evolution image text pairs. Use the harmfulness of evolution to help you make decision. Evolution: <Info> Requirement:
1. The output should be a word of either ’harmful’ or ’not harmful’. 2. meme is misogynous if Multi-modal unit consisting of an image and an embedded text that has the potential to cause harm to an individual, an organization, a community, or society. [Here is your expert response]
Table 11: Prompt design for final prediction of HarM.
Appendix E Related Work
Hateful Meme Detection.    Most works of hateful meme detection are based on pretrained vision-language models and fine-tune them on the hateful meme detection data. For example, Kiela et al. (2020) introduced the Hateful Memes Challenge and proposed a multimodal model that combines visual and textual features using a transformer-based architecture. Hate-CLIPper proposed a multimodal contrastive learning approach to improve the representation learning of memes by aligning the visual and textual modalities. These works demonstrate the effectiveness of using pretrained vision-language models. Differently, we discuss the potential of LMMs for hateful meme detection and introduce a knowledge-enhance LMM.
Large Multimodal Models.    With the release of the GPT-4V OpenAI (2023), many researchers focus on the development and application of LMMs in processing and understanding multimodal data. Discuss groundbreaking models like GPT-4V and LLAVA-1.5 Liu et al. (2023a), highlighting their capabilities in interpreting complex datasets that include a combination of text and images. Address the challenges these models face, such as sensitivity to noise and typographical errors, and the ongoing efforts to improve their accuracy and robustness in real-world applications. In this work, we adopt the LMMs to hateful meme detection, where set an LMM-based zero-shot hateful meme detection benchmark. Furthermore, we show that our Evolver further boosts the LMM’s performance.
More Cases Analysis.    To show the effectiveness of our method and align a better understanding of hatefulness in FHM dataset, we show more examples generated by Evolver and baseline in Fig. 4.
Figure 4: Example results of the Evolver (Ours) and the baseline model (MMICL).